You mix two sample of water. Sample A is 200g at 20oC. Sample B is 100g at 95oC. Calculate total change in entropy of the process. Assume that specific heat of water is 4190 J·g-1·K-1
1
Expert's answer
2019-06-10T09:41:48-0400

"m_1=0.2 kg" - mass of sample A

"m_2=0.1 kg" - mass of sample B

"T_1=20^0C+273=293K" - temperature of sample A

"T_2=95^0C+273=368K" - temperature of sample B

"c_{water}= 4190 Jkg^{-1}K^{-1}" //mistake in statements "g \to kg"

Lets agree the temperature of equilibrium "T_3"

According energy conversation law, heating of sample A equl to cooling sample B.


"c_{water}m_1(T_3-T_1) = c_{water}m_2(T_2-T_3) \implies"

"T_3=(m_2T_2+m_1T_1)\/(m_1+m_2)"


Using nubmber "T_3=318K"

"\Delta S_1=c_{water}m_1ln(T_3/T_1)" - entropy change for sample A

"\Delta S_2=c_{water}m_2ln(T_3/T_2)" - entropy change for sample B

"\Delta S=\Delta S_1+\Delta S_2" total entropy change

"\Delta S=7.4582 J/K"


Answer: "7.4582 J\/K"